# Transversal Structures in Graph Systems: A Survey

### v1

@@ 5 Random graph systems
In this section, we survey some results on the existence of a transversal structures in a system of random graphs in various settings. Let $0\leq p\leq 1$ and $G(n,p)$ be the binomial random graph. For graphs $G=(V,E)$ and $G^{\prime}=(V^{\prime},E^{\prime})$, we set $G\cap G^{\prime}=(V\cap V^{\prime},E\cap E^{\prime})$. Given an $n$-vertex graph $G$, we call $G(p):=G\cap G(n,p)$ the random subgraph of $G$. In particular, $G(p)=G(n,p)$ when $G$ is the complete graph $K_{n}.$ We say that $G(p)$ has a graph property $\mathcal{P}$ with high probability, or whp for brevity, if the probability that $G(p)$ has $\mathcal{P}$ approaches $1$ as $n$ tends to infinity. Sudakov [127] pointed out that random subgraph $G(p)$ can be used to measure the robustness of graph properties.
A natural question concerns the existence of transversal spanning structures in random graph systems. A graph property is called monotone increasing if it is closed under the addition of edges. For a monotone increasing graph property, the resilience quantifies the robustness in terms of the number of edges one must delete from a graph, locally or globally, in order to destroy the property. Ferber, Han and Mao [51] proved a resilient result for a system of $n$ graphs with order $n$ where each graph is an independent copy of $G(n,p)$ provided that $np=\omega(\log n)$.
@@
Similar result for transversal perfect matching was also obtained in [51]. Note that the above theorem implies that if $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq(1/2+\epsilon)n$, and =$\mathcal{F}=\{F_{1},\ldots,F_{n}\}$ where $F_{i}=G_{i}(p)$ for some $p\geq{c\log n}/n$, then whp $\mathcal{F}$ contains a transversal Hamilton cycle. Anastos and Chakraborti [11] improved this by removing the error term in the minimum degree condition with spread techniques.
@@
There exists an absolute $c$ such that the following holds. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq\frac{n}{2}$, and $p\geq\frac{c\log n}{n}$. Let =$\mathcal{F}=\{F_{1},\ldots,F_{n}\}$ where $F_{i}=G_{i}(p)$. Then, whp $\mathcal{F}$ contains a transversal Hamilton cycle.
In the above theorem, each graph in the given system intersect with the same random graph $G(n,p)$. They [11] also considered a system that consists of $n$ independent random subgraphs of a Dirac graph.
@@
Naturally, they [11] hoped for a combined generalization of the above two theorems.
@@
Does there exist a constant $c$ such that the following holds? Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a graph system of order $n$ with $\delta(\mathcal{G})\geq\frac{n}{2}$, and $p\geq\frac{c\log n}{n^{2}}$. For each $i\in[n]$, let $F_{i}=G_{i}(p)$ and let =$\mathcal{F}=\{F_{1},\ldots,F_{n}\}$. Then, whp $\mathcal{F}$ contains a transversal Hamilton cycle.
Anastos and Chakraborti [11] showed that Problem 5.4 holds when $n$ is odd, whereas it turns out to be false when $n$ is even.
Han, Hu and Yang [65] obtained a transversal robust version of the Hajnal–Szemerédi theorem [63].
@@
Let $r\geq 3$ be an integer and let $\gamma>0.$ There exists a constant $c=c(r,\gamma)$ such that for any sufficiently large $n\in r\mathbb{N}$ and $p\geq cn^{-\frac{2}{r}}(\log n)^{1/\binom{r}{2}}$ the following holds. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{\frac{n}{r}\binom{r}{2}}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq(1-\frac{1}{r}+\gamma)n$. Let =$\mathcal{F}=\{F_{1},\ldots,F_{\frac{n}{r}\binom{r}{2}}\}$ where $F_{i}=G_{i}(p)$. Then whp $\mathcal{F}$ contains a transversal $K_{r}$-factor.
@@ 6 Digraph systems
In contrast to Theorem 3.13, it is natural to seek an analogue of the transversal Hajnal–Szemerédi theorem [63] in the digraph systems. Cheng, Han, Wang and Wang [27] established the minimum out-degree and semi-degree condition for the existence of transversal tournament factors in digraph systems.
@@
    every system of $(n-1)$ tournaments with order $n$ contains a transversal directed Hamilton path;
@@
    every system of $(n-1)$ tournaments with order $n$ contains a transversal of every orientation of the Hamilton path;
@@
Babiński, Grzesik and Prorok [13] studied the transversal version of Mantel’s theorem in digraph systems and oriented graph systems. More precisely, they determined the minimum number of edges in each graph that ensure the existence of transversal transitive triangles or transversal directed triangles in digraph systems and oriented graph systems. Gerbner, Grzesik, Palmer and Prorok [58] determined the minimum number of edges of digraphs in a digraph system which guarantees the existence of any rainbow directed star.
@@
Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a system of digraphs on a common vertex set $V$ of sufficiently large order $n.$ If $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, then $\mathcal{D}$ contain a transversal directed Hamilton cycle.
DeBiasio and Molla [38] proved the the anti-directed Hamilton cycle (an orientation in which consecutive edges alternate direction) is guaranteed to appear if $\delta^{0}(D)\geq n/2+1$. DeBiasio, Kühn, Molla, Osthus and Taylor [37] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq n/2$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one. It is interesting to generalize the above two results to transversal version.
Keevash, Kühn and Osthus [77] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semi-degree at least $(3n-4)/8$ contains a directed Hamilton cycle. Considering a transversal version of this theorem would be interesting.
Question 6.5.
Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a system of oriented graphs with common vertex set $V$ of size $n.$ If $\delta^{0}(\mathcal{D})\geq\lceil\frac{3n-4}{8}\rceil$, does $\mathcal{D}$ contain a transversal directed Hamilton cycle?
@@ 7 Hypergraph systems
In this section, we will survey results in $k$-uniform hypergraph systems. For a set $V$, we denote by $\binom{V}{k}$ the set of subsets of size $k$ of $V$. A $k$-uniform hypergraph (or $k$-graph) ${H}=(V,E)$ consists of a vertex set $V$ and an edge set $E$ which is a family of $k$-element subsets of $V$, i.e., $E\subseteq\binom{V}{k}.$ For any $S\subseteq V$, the degree of $S$ in ${H}$, denoted by $\deg_{H}(S)$, is the number of edges containing $S.$ For any integer $s\geq 0$, define the minimum $s$-degree $\delta_{s}({H})$ to be $\min\{\deg_{H}(S):S\in\binom{V({H})}{s}\}$. Let =$\mathcal{H}=\{H_{1},H_{2},\dots,H_{m}\}$ be a system of not necessarily distinct $k$-graphs with common vertex set $V$. Define $\delta_{s}(\mathcal{H})=\min\{\delta_{s}(H_{i}):i\in[m]\}$.
@@
7.1 Transversal Hamilton cycle in hypergraph systems
We say that a $k$-graph ${C}$ is an $\ell$-cycle if there exists a cyclic ordering of the vertices of ${C}$ such that every edge of ${C}$ consists of $k$ consecutive vertices and every two consecutive edges (in the natural ordering of the edges) intersects in precisely $\ell$ vertices. We say that $C$ is a Hamilton $\ell$-cycle of a $k$-graph ${H}$ if $C$ is a spanning sub-$k$-graph of $H$. We call $C$ a tight Hamilton cycle if $\ell=k-1$, a loose Hamilton cycle if $\ell=1$. A Hamilton cycle of a $k$-graph $H$ is a cycle which is a spanning subhypergraph of $H$. Note that if a $k$-graph ${H}$ on $n$ vertices contains a Hamilton $\ell$-cycle then $(k-\ell)|n$, since every edge of the cycle contains exactly $k-\ell$ vertices which were not contained in the previous edge.
In 1999, Katona and Kierstead [76] proved a Dirac-type theorem: every $n$-vertex $k$-graph $H$ with $\delta_{k-1}({H})>(1-\frac{1}{2k})n+4-k-\frac{5}{2k}$ contains a tight Hamilton cycle. They anticipated that a much stronger result is true.
@@ 7.1 Transversal Hamilton cycle in hypergraph systems
Let $k\geq 3$, $\gamma>0$ and ${H}$ be an $n$-vertex $k$-graph, where $n$ is sufficiently large. If $\delta_{k-1}({H})\geq(\frac{1}{2}+\gamma)n$, then ${H}$ contains a tight Hamilton cycle. Furthermore, when $k=3$ it is enough to have $\delta_{2}({H})\geq\lfloor\frac{n}{2}\rfloor$.
However, Kühn and Osthus [96] and Zhao [135] noted that it is much more difficult to determine the minimum $d$-degree condition for tight Hamilton cycle for $d\in[k-2].$ The best general bound for such a problem was given by Lang and Sanhueza-Matamala [99], Polcyn, Reiher, Rödl and Schülke [119] independently.
@@
Let $k\geq 3$, $\gamma>0$ and ${H}$ be an $n$-vertex $k$-graph, where $n$ is sufficiently large. If $\delta_{k-2}({H})\geq(\frac{5}{9}+\gamma)\binom{n}{2}$, then ${H}$ contains a tight Hamilton cycle.
Let $11\leq d,\ell\leq k-$ 1. For $n\in(k-\ell)\mathbb{N}$, define $h_{d}^{\ell}(k,n)$ to be the smallest integer $h$ such that every $n$-vertex $k$-graph $H$ satisfying $\delta_{d}(H)\geq h$ contains a Hamilton $\ell$-cycle. Han and Zhao [66] proved that
@@
where $d\in[k-1]$ and $t=k-d$. In particular, $h_{k-3}^{k-1}(k,n)\geq(5/8+o(1))\binom{n}{3}$. Lang and Sanhueza-Matamala [99] proved $h_{d}^{k-1}(k,n)\leq 2^{-1/(k-d)}$ and they conjectured that the minimum $d$-degree threshold for $k$-uniform tight Hamilton cycles coincides with the lower bounds in (1).
Recently, Cheng, Han, Wang, Wang and Yang [28] and Tang, Wang, Wang and Yan [129] extended Theorem 7.2 and Theorem 7.3 to transversal version, respectively.
@@
For every $k\geq 3,\gamma>0$, there exists $n_{0}$ such that the following holds for $n\geq n_{0}.$ Let =$\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ be a $k$-graph system of order $n$.
@@
Cheng, Han, Wang, Wang and Yang [28] suspected that the threshold for transversal tight Hamilton cycles in a $k$-graph system is the same with the threshold conjectured for tight Hamilton cycles in a single $k$-graph (i.e., Conjecture 7.1) and provided the following conjecture.
@@
Suppose =$\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ is a $k$-graph system of order $n\,(n\geq k+1\geq 4)$ and $\delta_{k-1}(\mathcal{H})\geq\lfloor\frac{n-k+3}{2}\rfloor$, then $\mathcal{H}$ contains a transversal tight Hamilton cycle.
Motiveted by (1), Tang, Wang, Wang and Yan [129] proposed the following conjecture, which concerns the minimum $d$-degree condition for transversal tight Hamilton cycles.
@@
For every $k\geq d+1$, $\mu>0$, there exists $n_{0}$ such that the following holds for $n\geq n_{0}.$ Let =$\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ be a $k$-graph system of order $n$.
@@
Theorem 7.4 implies that if we consider $(k-1)$-degree and $(k-2)$-degree conditions, then tight Hamilton cycle is color-blind. Gupta, Hamann, Müyesser, Parczyk and Sgueglia [62] provided a unified approach to find color-blind graphs, which is general enough to recover many previous results and obtain novel transversal variants of several classical Dirac-type results. The precise statement of their main result is quite technical, we refer the reader to see [62, Theorem 2.6]. The following theorem lists some applications in their paper and they believed that their setting can capture even more families of hypergraphs.
@@
7.2 Transversal matching in hypergraph systems
A $t$-matching $M$ in a hypergraph $H$ is a subset of $E(H)$ consisting of $t$ pairwise disjoint edges, which is perfect if $V(M)=V(H)$. We denote by $\binom{[n]}{r}$ the set of subsets (also called edges) of size $r$ of $[n].$ For a set $X$, let $2^{X}$ be the set of all subsets of $X$. Let ${F}$ be a subfamily of $2^{X}$, denote by $\nu({F})$ its matching number, that is, the largest $m$ such that ${F}$ contains $m$ pairwise disjoint sets. A classical problem in extremal set theory is to determine the maximum size of ${F}$ with fixed $\nu({F})$. In 1965, Erdős [44] proposed the following conjecture, which is known as Erdős matching conjecture.

### v2

@@ 5 Random graph systems
In this section, we survey some results on the existence of transversal structures in a system of random graphs in various settings. Let $0\leq p\leq 1$ and $G(n,p)$ be the binomial random graph. For graphs $G=(V,E)$ and $G^{\prime}=(V^{\prime},E^{\prime})$, we set $G\cap G^{\prime}=(V\cap V^{\prime},E\cap E^{\prime})$. Given an $n$-vertex graph $G$, we call $G(p):=G\cap G(n,p)$ the random subgraph of $G$. In particular, $G(p)=G(n,p)$ when $G$ is the complete graph $K_{n}.$ We say that $G(p)$ has a graph property $\mathcal{P}$ with high probability, or whp for brevity, if the probability that $G(p)$ has $\mathcal{P}$ approaches $1$ as $n$ tends to infinity. Sudakov [134] pointed out that $G(p)$ can be used to measure the robustness of graph properties.
A natural question concerns the existence of transversal spanning structures in random graph systems. A graph property is called monotone increasing if it is closed under the addition of edges. For a monotone increasing graph property, the resilience quantifies the robustness in terms of the number of edges one must delete from a graph, locally or globally, in order to destroy the property. Ferber, Han and Mao [53] proved a resilient result for a system of $n$ graphs with order $n$ where each graph is an independent copy of $G(n,p)$ provided that $np=\omega(\log n)$, and they also noted that the bound for $p$ is probably not tight.
@@
A similar result for transversal perfect matchings was also obtained in [53]. Note that the above theorem implies that if $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq(1/2+\epsilon)n$, and $\mathcal{F}=\{F_{1},\ldots,F_{n}\}$ where $F_{i}=G_{i}(p)$ for some $p\geq{c\log n}/n$, then whp $\mathcal{F}$ contains a transversal Hamilton cycle. Anastos and Chakraborti [11] improved this by removing the error term in the minimum degree condition with spread techniques.
@@
There exists an absolute $c$ such that the following holds. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq\frac{n}{2}$, and $p\geq\frac{c\log n}{n}$. Let $\mathcal{F}=\{F_{1},\ldots,F_{n}\}$ where $F_{i}=G_{i}(p)$. Then, whp $\mathcal{F}$ contains a transversal Hamilton cycle.
In the above theorem, each graph in the given system intersects with the same random graph $G(n,p)$. They [11] also considered a system that consists of $n$ independent random subgraphs of a Dirac graph.
@@
In Theorem 5.2 and Theorem 5.3, the lower bound on $p$ is optimal up to a constant factor $c$. Naturally, they [11] hoped for a combined generalization of the above two theorems.
@@
Does there exist a constant $c$ such that the following holds? Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a graph system of order $n$ with $\delta(\mathcal{G})\geq\frac{n}{2}$, and $p\geq\frac{c\log n}{n^{2}}$. For each $i\in[n]$, let $F_{i}=G_{i}(p)$ and let $\mathcal{F}=\{F_{1},\ldots,F_{n}\}$. Then, whp $\mathcal{F}$ contains a transversal Hamilton cycle.
In the same paper, they [11] showed that Problem 5.4 has a positive answer when $n$ is odd, whereas it turns out to be false when $n$ is even.
Han, Hu and Yang [68] obtained a transversal robust version of the Hajnal–Szemerédi theorem [66], where the bound for $p$ is not tight.
@@
Let $r\geq 3$ be an integer and let $\gamma>0.$ There exists a constant $c=c(r,\gamma)$ such that for any sufficiently large $n\in r\mathbb{N}$ and $p\geq cn^{-\frac{2}{r}}(\log n)^{1/\binom{r}{2}}$ the following holds. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{\frac{n}{r}\binom{r}{2}}\}$ is a graph system of order $n$ with $\delta(\mathcal{G})\geq(1-\frac{1}{r}+\gamma)n$. Let $\mathcal{F}=\{F_{1},\ldots,F_{\frac{n}{r}\binom{r}{2}}\}$ where $F_{i}=G_{i}(p)$. Then whp $\mathcal{F}$ contains a transversal $K_{r}$-factor.
@@ 6 Digraph systems
In light of Theorem 3.13, it is natural to seek an analogue of the transversal Hajnal–Szemerédi theorem [66] in digraph systems. Cheng, Han, Wang and Wang [28] established the minimum out-degree and semi-degree condition for the existence of transversal tournament factors in digraph systems.
@@
    every system of (n1)(n-1) tournaments with order $n$ contains a transversal directed Hamilton path;
@@
    every system of (n1)(n-1) tournaments with order $n$ contains a transversal of every orientation of the Hamilton path;
@@
Babiński, Grzesik and Prorok [13] studied the transversal version of Mantel’s theorem in digraph systems and oriented graph systems. More precisely, they determined the minimum number of edges in each graph that ensures the existence of transversal transitive triangles or transversal directed triangles in digraph systems and oriented graph systems. Gerbner, Grzesik, Palmer and Prorok [60] determined the minimum number of edges required in each digraph of a digraph system that guarantees the existence of any rainbow directed star.
@@
Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a system of digraphs on a common vertex set $V$ of sufficiently large order $n.$ If $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
DeBiasio and Molla [40] proved the anti-directed Hamilton cycle (an orientation in which consecutive edges alternate direction) is guaranteed to appear if $\delta^{0}(D)\geq n/2+1$. DeBiasio, Kühn, Molla, Osthus and Taylor [39] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq n/2$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one. It would be interesting to generalize the above two results to the transversal setting.
Keevash, Kühn and Osthus [83] proved the oriented version of Dirac’s theorem, which states that for sufficiently large $n$, every $n$-vertex digraph with minimum semi-degree at least (3n4)/8(3n-4)/8 contains a directed Hamilton cycle. Considering a transversal version of this theorem would be interesting.
Problem 6.5 ([30]).
Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a system of oriented graphs with common vertex set $V$ of size $n.$ For sufficiently large $n$, if $\delta^{0}(\mathcal{D})\geq\lceil\frac{3n-4}{8}\rceil$, does $\mathcal{D}$ contain a transversal directed Hamilton cycle?
@@ 7 Hypergraph systems
In this section, we will survey results in $k$-uniform hypergraph systems. For a set $V$, we denote by $\binom{V}{k}$ the set of subsets of size $k$ of $V$. A $k$-uniform hypergraph (or $k$-graph) H=(V,E){H}=(V,E) consists of a vertex set $V$ and an edge set $E$ which is a family of $k$-element subsets of $V$, i.e., $E\subseteq\binom{V}{k}$. For any $S\subseteq V$, the degree of $S$ in H{H}, denoted by $\deg_{H}(S)$, is the number of edges containing $S.$ For any integer $s\geq 0$, define the minimum $s$-degree $\delta_{s}({H})$ to be $\min\{\deg_{H}(S):S\in\binom{V({H})}{s}\}$. Let $\mathcal{H}=\{H_{1},H_{2},\dots,H_{m}\}$ be a system of not necessarily distinct $k$-graphs with common vertex set $V$. Define $\delta_{s}(\mathcal{H})=\min\{\delta_{s}(H_{i}):i\in[m]\}$.
@@
7.1 Hamilton cycle
We say that a $k$-graph C{C} is an $\ell$-cycle if there exists a cyclic ordering of the vertices of C{C} such that every edge of C{C} consists of $k$ consecutive vertices and every two consecutive edges (in the natural ordering of the edges) intersects in precisely $\ell$ vertices. We say that $C$ is a Hamilton $\ell$-cycle of a $k$-graph H{H} if $C$ is a spanning sub-$k$-graph of $H$. We call $C$ a tight Hamilton cycle if $\ell=k-1$, a loose Hamilton cycle if $\ell=1$. A Hamilton cycle of a $k$-graph $H$ is a cycle which is a spanning subhypergraph of $H$. Note that if a $k$-graph H{H} on $n$ vertices contains a Hamilton $\ell$-cycle then $(k-\ell)|n$, since every edge of the cycle contains exactly $k$-$\ell$ vertices which were not contained in the previous edge.
In 1999, Katona and Kierstead [82] proved a Dirac-type theorem in hypergraphs: every $n$-vertex $k$-graph $H$ with $\delta_{k-1}({H})>(1-\frac{1}{2k})n+4-k-\frac{5}{2k}$ contains a tight Hamilton cycle. They anticipated that a much stronger result is true.
@@ 7.1 Hamilton cycle
Let $k\geq 3$, $\gamma>0$ and H{H} be an $n$-vertex $k$-graph, where $n$ is sufficiently large. If $\delta_{k-1}({H})\geq(\frac{1}{2}+\gamma)n$, then H{H} contains a tight Hamilton cycle. Furthermore, when $k=3$ it is enough to have $\delta_{2}({H})\geq\lfloor\frac{n}{2}\rfloor$.
However, Kühn and Osthus [102] and Zhao [144] noted that it is much more difficult to determine the minimum $d$-degree condition for tight Hamilton cycle for $d\in[k-2].$ The best general bound for such a problem was given by Lang and Sanhueza-Matamala [106], and Polcyn, Reiher, Rödl and Schülke [126] independently.
@@
Let $k\geq 3$, $\gamma>0$ and H{H} be an $n$-vertex $k$-graph, where $n$ is sufficiently large. If $\delta_{k-2}({H})\geq(\frac{5}{9}+\gamma)\binom{n}{2}$, then H{H} contains a tight Hamilton cycle.
Let $1\leq d,\ell\leq k-1$. For $n\in(k-\ell)\mathbb{N}$, define $h_{d}^{\ell}(k,n)$ to be the smallest integer $h$ such that every $n$-vertex $k$-graph $H$ satisfying $\delta_{d}(H)\geq h$ contains a Hamilton $\ell$-cycle. Han and Zhao [69] proved that
@@
where $d\in[k-1]$ and t=kdt=k-d. In particular, $h_{k-3}^{k-1}(k,n)\geq(5/8+o(1))\binom{n}{3}$. Lang and Sanhueza-Matamala [106] proved $h_{d}^{k-1}(k,n)\leq 2^{-1/(k-d)}$ and they conjectured that the minimum $d$-degree threshold for $k$-uniform tight Hamilton cycles coincides with the lower bounds in (1).
Recently, Cheng, Han, Wang, Wang and Yang [29] and Tang, Wang, Wang and Yan [137] extended Theorem 7.2 and Theorem 7.3 to transversal versions, respectively.
@@
For every $k\geq 3,\gamma>0$, there exists $n_{0}$ such that the following holds for $n\geq n_{0}.$ Let $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ be a $k$-graph system of order $n$.
@@
Cheng, Han, Wang, Wang and Yang [29] suspected that the threshold for transversal tight Hamilton cycles in a $k$-graph system is the same as the threshold conjectured for tight Hamilton cycles in a single $k$-graph (i.e., Conjecture 7.1) and provided the following conjecture.
@@
Suppose $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ is a $k$-graph system of order $n\,(n\geq k+1\geq 4)$ and $\delta_{k-1}(\mathcal{H})\geq\lfloor\frac{n-k+3}{2}\rfloor$, then $\mathcal{H}$ contains a transversal tight Hamilton cycle.
Motivated by (1), Tang, Wang, Wang and Yan [137] proposed the following conjecture, which concerns the minimum $d$-degree condition for transversal tight Hamilton cycles.
@@
For every $k\geq d+1$, $\mu>0$, there exists an integer $n_{0}$ such that the following holds for $n\geq n_{0}.$ Let $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ be a $k$-graph system of order $n$.
@@
Theorem 7.4 implies that if we consider (k1)(k-1)-degree and (k2)(k-2)-degree conditions, then tight Hamilton cycles are color-blind. Gupta, Hamann, Müyesser, Parczyk and Sgueglia [65] provided a unified approach to find color-blind graphs, which is general enough to recover many previous results and obtain novel transversal variants of several classical Dirac-type results. The precise statement of their main result is quite technical, we refer the reader to see [65, Theorem 2.6]. The following theorem lists some applications in their paper and they believed that their setting can capture even more families of hypergraphs.
@@
7.2 Matching and factor
A $t$-matching $M$ in a hypergraph $H$ is a subset of $E(H)$ consisting of $t$ pairwise disjoint edges, which is perfect if $V(M)=V(H)$. We denote by $\binom{[n]}{r}$ the set of subsets (also called edges) of size $r$ of $[n].$ For a set $X$, let $2^{X}$ be the set of all subsets of $X$. Let F{F} be a subfamily of $2^{X}$, denote by $\nu({F})$ its matching number, that is, the largest $m$ such that F{F} contains $m$ pairwise disjoint sets. A classical problem in extremal set theory is to determine the maximum size of F{F} with fixed $\nu({F})$. In 1965, Erdős [46] proposed the following conjecture, which is known as Erdős matching conjecture.